What minimum force applied at 8 inches axle of 2 feet wheel is necessary to raise the wheel over an obstacle of 10 inches height? Take r as radius and W as its weight.
1
Expert's answer
2019-10-14T09:31:07-0400
"r=2\ ft=0.6096\ m, h=10\ in=0.254\ m, m=1.17\ kg"

"a^2+(r-h)^2=r^2\to a^2+(0.6096-0.254)^2=0.6096^2"

"a=0.495\ m"

Moments of two forces will be equal:


"F(r-h)=mga"

"F(0.6096-0.254)=(1.17)(9.8)(0.495)"

"F=16.0\ N"